# TAMMs: Temporal-Aware Multimodal Model for Satellite Image Change Understanding and Forecasting

### v2

@@ 1 Introduction
Understanding temporal change and forecasting future scenes from Satellite Image Time Series (SITS) is fundamental to critical applications Fu et al. (2024); Lara-Alvarez et al. (2024), yet it poses a profound challenge: predicting the future requires a deep understanding of the past change. While foundation models like Multimodal Large Language Models (MLLMs) Wu et al. (2024); Bai et al. (2025) and diffusion-based generators Sastry et al. (2024); Liu et al. (2025a) have demonstrated remarkable capabilities, their application to this domain remains fragmented and fundamentally limited. As illustrated in Fig. 1, the field is split into disjointed paradigms: static image captioning that ignores temporal context Bazi et al. (2024), and temporal description methods that lack semantic depth of MLLMs Peng et al. (2024); similarly, generative approaches are divided between single-image synthesis that cannot model evolution Liu et al. (2025a), and temporal forecasting struggles with long-term consistency because fails to grasp semantic drivers of change Khanna et al. (2024). This division reflects a scientific gap —— the absence of a unified framework of temporal-aware multimodal model for satellite image change understanding and forecasting named TAMMs, which can reason about why changes occur and use that reasoning to generate what will happen next.
How can model reason about why these temporal changes occur? Existing Temporal Change Description methods fuse temporal multi-graph information through simple interactions, and their long-range temporal reasoning capabilities are limited Peng et al. (2024). To bridge this gap, we introduce novel, parameter-efficient approach to awaken the latent temporal reasoning abilities of frozen MLLM through Temporal Adaptation Modules (TAM), designed to both inject explicit awareness of arbitrary time intervals and guide the model’s focus onto salient change narratives, enhancing its performance in long-range temporal change reasoning.
@@ 2 Task Definition
Our work introduces a unified framework to address two coupled, yet historically disjointed, tasks central to Satellite Image Time Series (SITS) analysis. We argue that solving them synergistically is key to unlocking more reliable, reasoning-based forecasting.
@@ 4.1.1 Datasets
We evaluate our model’s performance on two tasks. For assessing the textual descriptions summarizing temporal changes, we utilize standard natural language generation metrics: BLEU-4 Papineni et al. (2002), METEOR Lavie & Agarwal (2007), ROUGE-L Lin (2004), and CIDEr-D Vedantam et al. (2015), which collectively measure aspects like n-gram precision, word alignment considering synonyms, longest common subsequence overlap, and consensus among reference descriptions. To evaluate predicted images, besides standard metrics PSNR Zhang et al. (2020), SSIM Wang et al. (2004), and LPIPS Zhang et al. (2018), we further introduce TCS metric for evaluating temporal image generation, which has detailed in Section 2.3.
@@ 4.1.3 Implementation Details
Our framework is conducted on two NVIDIA A6000 GPUs. The MLLM backbone is a frozen DeepSeek-VL2 Wu et al. (2024), while the generative component adapts the U-Net architecture from DiffusionSat Khanna et al. (2024). A critical component of our methodology is a multi-stage training strategy for the generative module. This approach is designed to first establish robust spatio-temporal structural control before integrating the MLLM’s high-level semantic guidance. This decoupling prevents catastrophic forgetting and ensures a stable fusion of the low-level structural priors and high-level semantic reasoning. A detailed breakdown of our three-stage training protocol, including specific learning rates and optimizer schedules for each stage, is provided in Appendix A.8.2. Comprehensive architectural details and a full list of hyperparameters can also be found in Table 4.
@@ 4.2.2 Quantitative Results
In temporal change description, TAMMs outperforms prior methods such as SITSCC (0.2669 vs. 0.2122 BLEU-4, 0.3312 vs. 0.2961 METEOR) and DeepSeek-VL2 (0.9030 vs. 0.7362 CIDEr-D), highlighting its enhanced temporal reasoning and generation fidelity. For future satellite image prediction, TAMMs achieves the highest PSNR (12.07), competitive SSIM (0.1831 vs. 0.2098 for MCVD), and the best LPIPS (0.4931). While the improvements in standard quality metrics (PSNR, SSIM, LPIPS) are modest, TAMMs demonstrates a substantial advantage in TCS, achieving 0.9690, compared to 0.7624 for DiffusionSat and only 0.2170 for GeoSynth-Canny. This TCS margin indicates TAMMs’ superior ability to produce temporally coherent futures grounded in prior dynamics.
@@ A.4 Contextual Temporal Prompt
A.5 Training Strategy for the Generative Module
@@
A.5.1 Stage 1: Structural Control Pre-training
@@
A.5.2 Stage 2: Semantic and Temporal Refinement
@@
A.6 Dataset Details
A.6.1 Dataset Generating Prompts
@@
A.6.2 Dataset Distribution
@@
A.7 Concrete Generated Text Length Analysis
@@
A.8 Training Details
A.8.1 Model Architecture and Hyperparameters
@@
A.8.2 Multi-Stage Training Strategy
@@
The initial stage focuses on learning the foundational spatio-temporal dynamics. We freeze the U-Net and MLLM, training only the 3D Control Block weights ($\theta_{st}$). This allows the model to learn to extract structural priors from the image sequence alone, without semantic influence. This stage was trained for approximately 50k steps using the AdamW optimizer Kingma & Ba (2014) with a constant learning rate of 2e-5.

### v3

# TAMMs: Change Understanding and Forecasting in Satellite Image Time Series with Temporal-Aware Multimodal Models

@@ 1 Introduction
Understanding temporal change and forecasting future scenes from Satellite Image Time Series (SITS) is fundamental to critical applications Fu et al. (2024); Lara-Alvarez et al. (2024), yet it poses a profound challenge: predicting the future requires a deep understanding of the past change. While foundation models like Multimodal Large Language Models (MLLMs) Wu et al. (2024); Bai et al. (2025) and diffusion-based generators Sastry et al. (2024); Liu et al. (2025a) have demonstrated remarkable capabilities, their application to this domain remains fragmented and fundamentally limited. As illustrated in Fig. 1, the field is split into disjointed paradigms: static image captioning that ignores temporal context Bazi et al. (2024), and temporal description methods that lack semantic depth of MLLMs Peng et al. (2024); similarly, generative approaches are divided between single-image synthesis that cannot model evolution Liu et al. (2025a), and temporal forecasting struggles with long-term consistency because they fail to grasp semantic drivers of change Khanna et al. (2024). This division reflects a scientific gap —— the absence of a unified framework of temporal-aware multimodal model for satellite image change understanding and forecasting. To address this, We present TAMMs which can reason about why changes occur and use that reasoning to generate what will happen next.
How can model reasons about why these temporal changes occur? Existing Temporal Change Description methods fuse temporal multi-graph information through simple interactions, and their long-range temporal reasoning capabilities are limited Peng et al. (2024). To bridge this gap, we introduce novel, parameter-efficient approach to awaken the latent temporal reasoning abilities of frozen MLLM through Temporal Adaptation Modules (TAM), designed to both inject explicit awareness of arbitrary time intervals and guide the model’s focus onto salient change narratives, enhancing its performance in long-range temporal change reasoning.
@@ 2 Task Definition
Our work introduces a unified framework to address two coupled, yet historically disjointed, tasks central to SITS analysis. We argue that solving them synergistically is key to unlocking more reliable, reasoning-based forecasting.
@@ 4.1.1 Datasets
We evaluate our model’s performance on two tasks. For assessing the textual descriptions summarizing temporal changes, we utilize standard natural language generation metrics: BLEU-4 Papineni et al. (2002), METEOR Lavie and Agarwal (2007), ROUGE-L Lin (2004), and CIDEr-D Vedantam et al. (2015), which collectively measure aspects like n-gram precision, word alignment considering synonyms, longest common subsequence overlap, and consensus among reference descriptions. To evaluate predicted images, besides standard metrics PSNR Zhang et al. (2020), SSIM Wang et al. (2004), and LPIPS Zhang et al. (2018), we further introduce TCS metric for evaluating temporal image generation, which has detailed in Section 2.3.
@@ 4.1.3 Implementation Details
Our framework is conducted on two NVIDIA A6000 GPUs. The MLLM backbone is a frozen DeepSeek-VL2 Wu et al. (2024) due to its MoE architecture and open-source accessibility, while the generative component adapts the U-Net architecture from DiffusionSat Khanna et al. (2024) which is based on Stable Diffusion 2-1. We initially considered native unified models (Chameleon Team (2024), Janus-Pro Chen et al. (2025)) which could theoretically leverage cross-task information more effectively, but preliminary experiments showed significant performance gaps compared to specialized models, making them unsuitable for demanding remote sensing tasks.
A critical component of our methodology is a multi-stage training strategy for the generative module. This approach is designed to first establish robust spatio-temporal structural control before integrating the MLLM’s high-level semantic guidance. This decoupling prevents catastrophic forgetting and ensures a stable fusion of the low-level structural priors and high-level semantic reasoning. A detailed breakdown of our three-stage training protocol, including specific learning rates and optimizer schedules for each stage, is provided in Appendix A.9.2. Comprehensive architectural details and a full list of hyperparameters can also be found in Table 5.
@@ 4.2.2 Quantitative Results
In temporal change description, TAMMs outperforms prior methods such as SITSCC (0.2669 vs. 0.2122 BLEU-4, 0.3312 vs. 0.2961 METEOR) and DeepSeek-VL2 (0.9030 vs. 0.7362 CIDEr-D), highlighting its enhanced temporal reasoning and generation fidelity. For future satellite image prediction, TAMMs achieves the highest PSNR (12.07), competitive SSIM (0.1831 vs. 0.2098 for MCVD), and the best LPIPS (0.4931).
We notice the improvements in standard quality metrics (SSIM) are modest, this pattern is consistent with recent diffusion-based forecasting literature, where models that better preserve high-frequency and structural details often trade a small amount of PSNR/SSIM for more realistic and temporally consistent predictions Khanna et al. (2024). In our setting, such perceptual–temporal fidelity (captured by LPIPS and especially TCS) is more critical than maximizing distortion-based scores, because downstream change analysis depends on plausible evolution rather than exact pixel-wise reconstruction. TAMMs demonstrates a substantial advantage in TCS, achieving 0.9690, compared to 0.7624 for DiffusionSat and only 0.2170 for GeoSynth-Canny. This TCS margin indicates TAMMs’ superior ability to produce temporally coherent futures grounded in prior dynamics.
@@ 6 Conclusion and Future Work
Acknowledgements
This work was supported by the Foundation Program of the Key Laboratory of Science and Technology on Complex Electronic System Simulation under Grant 614201001032203. We are grateful to our colleagues and research experts in remote sensing and aerospace for their constructive discussions and guidance throughout this work.
@@ A.4 Contextual Temporal Prompt
A.5 Prompt Robustness Analysis Details
To evaluate the robustness of our Contextual Temporal Prompting (CTP), we conducted an additional ablation study using three variants of prompts with varying levels of detail to test the stability of our model:
  • $\mathcal{P}_{Short}$: “Describe the changes in the satellite images.” (Minimal instruction).
  • $\mathcal{P}_{Ours}$ (Default): The structured prompt used in the paper (see Appendix A.4).
  • $\mathcal{P}_{Verbose}$: A complex prompt adding approximately 50 words of context about specific land-cover types.
The full text for the $\mathcal{P}_{Verbose}$ prompt is as follows:
  • <image>...<image> Scene: [Scene Description]. Analyze the provided satellite image time series with particular attention to diverse land-cover categories, including urban infrastructure, agricultural vegetation patterns, and hydrological features. Please distinguish between transient seasonal effects (e.g., phenological changes) and permanent structural developments. Based on this context, describe specific changes between these time-series remote sensing images in a single paragraph. Focus on concrete changes to structures, landscape, or development with precise location details.
The quantitative results are presented in Table 3. As shown, the performance variance is minimal (TCS fluctuates <0.015<0.015). This demonstrates that TAM is robust: the learnable visual alignment plays the primary role, while the prompt serves as a high-level task trigger.
Table 3: Prompt Robustness Analysis.
Prompt Type TCS BLEU-4 CIDEr-D
$\mathcal{P}_{Short}$ 0.9580 0.2610 0.8910
$\mathcal{P}_{Ours}$ (Default) 0.9690 0.2669 0.9030
$\mathcal{P}_{Verbose}$ 0.9615 0.2675 0.8980
A.6 Training Strategy for the Generative Module
@@
A.6.1 Stage 1: Structural Control Pre-training
@@
A.6.2 Stage 2: Semantic and Temporal Refinement
@@
A.7 Dataset Details
A.7.1 Dataset Generating Prompts
@@
Figure 7: Samples of Satellite Image Time Series captions TAMMs used.
-1em
@@
A.7.2 Scene Captions Samples
Figure 7 presents samples of Satellite Image Time Series captions generated by Qwen2.5-VL Bai et al. (2025).
A.7.3 Dataset Distribution
@@
A.8 Concrete Generated Text Length Analysis
@@
A.9 Training Details
A.9.1 Model Architecture and Hyperparameters
@@
A.9.2 Multi-Stage Training Strategy
@@
The initial stage focuses on learning the foundational spatio-temporal dynamics. We freeze the U-Net and MLLM, training only the 3D Control Block weights ($\theta_{st}$). This allows the model to learn to extract structural priors from the image sequence alone, without semantic influence. This stage was trained for approximately 50k steps using the AdamW optimizer Kingma and Ba (2014) with a constant learning rate of 2e-5.